# Machine Learning-based Unfolding for Cross Section Measurements in the Presence of Nuisance Parameters

### v1

@@ 3.2 More details on nuisance parameters update
During each iteration, the nuisance parameters are updated according to Eq. (23). Note that in the equation, $w(y,x,\theta^{(k)}),\nu^{(k)}(x)$ were already computed in the previous iteration, $\frac{p(y)}{\tilde{q}^{(k)}(y)}$ is the density ratio being estimated by the step-1 classifier, and $\log p_{0}(\theta)$ is known. The only unknown term is $\log[w(y,x,\theta)]$. Combined with the other quantities, this term is integrated with respect to the joint density $q(x,y)$, which can be approximated by a sample average over the MC data. Therefore, the only challenge lies in the estimation of the function $w(y,x,\theta)$. Recall that $w(y,x,\theta)$ is defined as the ratio of two conditional densities $\frac{p(y|x,\theta)}{q(y|x)}$. Chan2023 proposed to learn $w(y,x,\theta)$ by factorizing it into the product of two density ratios
@@
Let $X_{i}\sim\mathcal{P}_{X},\theta_{i}\sim\mathcal{P}_{\theta},Y_{i}\sim p(\cdot|X_{i},\theta_{i})$, and $X_{i}^{\prime}\sim\mathcal{Q}_{X},\theta_{i}^{\prime}\sim\mathcal{Q}_{\theta},Y_{i}^{\prime}\sim q(\cdot|X_{i})$. Let $f_{1}:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow[0,1]$ be the Bayes optimal classifier discriminating dataset $\mathcal{D}_{1}=\{X_{i},Y_{i},\theta_{i}\}$ from $\mathcal{D}_{2}=\{X^{\prime}_{i},Y^{\prime}_{i},\theta_{i}^{\prime}\}$. Let $f_{2}:\mathcal{X}\times\Theta\rightarrow[0,1]$ be the Bayes optimal classifier discriminating dataset $\tilde{\mathcal{D}}_{2}=\{X^{\prime}_{i},\theta^{\prime}_{i}\}$ from $\tilde{\mathcal{D}}_{1}=\{X_{i},\theta_{i}\}$. Then the output of the classifiers satisfy:
@@
The procedure outlined in Proposition 3 was also used in Chan2023, although a formal proof was not provided. In this setting, the auxiliary variable $\theta_{i}^{\prime}$ does not influence the synthetic data $X_{i}^{\prime},Y_{i}^{\prime}$; rather, it functions solely as a supporting variable for classifier training. While the distributions of $\theta_{i}$ and $\theta_{i}^{\prime}$ may differ, as allowed by the proposition, in practice, it might be preferable to set $q_{\theta}=p_{\theta}$ to avoid potential numerical instability. A convenient choice is to use a uniform distribution over the parameter space. Similarly, although the proposition specifies $X_{i}\sim\mathcal{P}_{X},X_{i}^{\prime}\sim\mathcal{Q}_{X}$, there is no restriction against using the same distribution for both. Therefore, a practical procedure can be summarized as follows:
@@
    Sample $X_{i}^{\prime},X_{i}$ from the particle-level MC distribution $q(x)$.
@@ 4.2 Neural network architecture and training
The neural network classifier for estimating density ratios during POF iteration is implemented in TensorFlow and Keras (tensorflow; keras). The network contains three hidden layers with 50 nodes per layer and employs the ReLU activation function. The output layer consists of a single node with a sigmoid activation function. Training is performed with the Adam optimizer (kingma2017adammethodstochasticoptimization) with learning rate $\eta=0.001$ using a weighted binary cross-entropy loss. The model is trained for up to 20 epochs with a batch size of 10,000, and early stopping with a patience of 3 epochs is applied, that is, training stops if the validation loss does not improve for 3 consecutive epochs. None of the hyperparameters were optimized.
Figure 4: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 2. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
The neural network classifiers for estimating the $w$ function ($w(y,x,\theta)=p(y|x,\theta)/q(y|x)$) are implemented in PyTorch (paszke2019pytorch). [^2] The reason we used two different frameworks is that the original implementation of OmniFold was in TensorFlow/Keras, while the codebase for learning $w$ function from Chan2023 was in PyTorch. There is no technical reason preventing from using a single framework. The same architecture is employed for both classifiers in Proposition 3. The network contains three hidden layers with 50 nodes per layer and employs the ReLU activation function. Batch normalization is applied after each hidden layer, and a dropout layer with a rate of 0.1 is added after the second hidden layer. The output layer consists of a single node with a sigmoid activation function. Training uses the Adam optimizer with learning rate $\eta=0.001$ and the weighted binary cross-entropy loss. This classifier is trained for up to 1000 epochs with a batch size of 10,000, and early stopping with patience 10 is used. In addition, we train an ensemble of 10 networks with bootstrap resampling to reduce the variance of the estimated w function. The range of nuisance parameter $\theta$ used in training is set to be $[0.5,2.0]$.
@@ 4.3 Results
Figure 7: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 5. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
Moreover, Figure 4 shows the evolution of the nuisance parameter $\theta$ and the goodness-of-fit statistic for the step-1 classifier, defined in Eq. (25). The top plot shows that the nuisance parameter converges to the true value within a few iterations, regardless of the initial value. The bottom plot shows the goodness-of-fit statistic converging to 1 for all initial values, indicating that the reweighted distribution $\tilde{q}(y)$ is close to the target distribution $p(y)$. As discussed in Section 3.1, the results shown in Figure 2-3 correspond to the solution with the highest goodness-of-fit statistic, although in this case study all solutions give equally good fits after a few iterations.
@@ 5.1 Dataset
To demonstrate this setup in practice, we use high-fidelity simulations from the CMS Experiment at the Large Hadron Collider. In particular, the data were generated with Pythia 6.426 (Sjostrand:2006za) using the Z2 tune (Chatrchyan:2011id) and interfaced with a Geant4-based (Agostinelli:2002hh; 1610988; Allison:2016lfl) detailed detector simulation of the CMS experiment (Chatrchyan:2008aa). This dataset comes from the CMS Open Data Portal (CMS:QCDsim1000-1400; CMS:QCDsim1400-1800; CMS:QCDsim1800) and is processed into an MIT Open Data format (komiske_patrick_2019_3341502; komiske_patrick_2019_3341770; komiske_patrick_2019_3341772; Komiske:2019jim). We use the events as both “simulation” and “data” in order to have a known target for testing.
The datasets from this collection are sorted by the parton-level hard-scattering scale $\hat{p}_{T}$ from Pythia, which is in general different from the jet-level transverse momentum $p_{T}$ we are interested in studying. For simplicity, we consider one slice of the collection with 600 GeV $\hat{p}_{T}<$ 800 GeV. This slice corresponds to sufficiently high momentum jets that effects from the triggering system are not relevant. Particles (at truth level) or particle flow candidates (at reconstructed/detector level) are used as inputs to jet clustering, implemented using FastJet 3.2.1 (Cacciari:2011ma; Cacciari:2005hq) and the anti-$k_{t}$ algorithm (Cacciari:2008gp) with radius parameter $R=0.5$. The total number of events is 43,892. We consider the two highest $p_{T}$ jets at both particle (truth) and detector (reconstructed) levels, denoted as $p_{T,1}^{\text{truth}},p_{T,2}^{\text{truth}},p_{T,1}^{\text{reco}},p_{T,2}^{\text{reco}}$. Then our observables are
@@ 5.3 Results
Moreover, Figure 10 shows that the nuisance parameter does not necessarily converge to a good estimate. In particular, when initialized at $\theta^{(0)}=1.0$ or $1.1$, the estimated value stabilizes around $\hat{\theta}\approx 1.35$. In these cases, the corresponding goodness-of-fit statistic also fails to approach 1, indicating that the reweighted distribution $\tilde{q}(y)$ does not adequately match the observed distribution $p(y)$. This behavior suggests that the algorithm may converge to a local, rather than global, maximum of the likelihood if the starting value $\theta^{(0)}$ is far from the optimal value. The results reported in Figures 8-9 are obtained by selecting, among all runs, the solution achieving the highest goodness-of-fit statistic, thereby avoiding this issue. Additional experiments with different jet energy resolution are provided in the supplementary material. These experiments yield consistent results: POF reliably recovers the true particle-level distribution, whereas OF fails in the presence of a misspecified nuisance parameter. However, the choice of the initial nuisance parameter remains important for POF, as starting too far from the true value can lead to suboptimal estimates. For both OF and POF, the reweighted detector-level distributions may also fail to perfectly match the experimental data, particularly when the discrepancy between the true $\theta$ and the nominal $\theta$ used in simulation is large. These findings highlight the importance of employing multiple initializations and selecting the best solution based on the goodness-of-fit statistic.
Figure 10: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 8. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@ 6 Discussion
Another important direction for future work is uncertainty quantification of the unfolding results. The current POF algorithm does not provide uncertainty estimates for either the unfolded distribution or the nuisance parameters. Even in the original OmniFold algorithm, it is unclear how to propagate the uncertainty in the classifier-based density ratio estimates to the final unfolded result. One possible (computationally expensive) solution is to use bootstrap resampling (Canelli2025UnbinnedUnfolding) to estimate the uncertainty of the unfolded distribution and nuisance parameters. However, rigorous analysis of uncertainty quantification still remains an open problem in this setting.
@@ A.1 Gaussian Data
    Figure 13: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 11. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 16: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 14. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 19: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 17. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 22: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 20. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@ A.2 CMS Open Data
    Figure 25: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 23. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 28: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 26. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 31: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 29. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 34: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the results shown in Figure 32. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.

### v2

@@ 3.2 More details on nuisance parameters update
During each iteration, the nuisance parameters are updated according to Eq. (23). Note that in the equation, $w(y,x,\theta^{(k)}),\nu^{(k)}(x)$ were already computed in the previous iteration, $\frac{p(y)}{\tilde{q}^{(k)}(y)}$ is the density ratio being estimated by the step-1 classifier, and $\log p_{0}(\theta)$ is known. The only unknown term is $\log[w(y,x,\theta)]$. Combined with the other quantities, this term is integrated with respect to the joint density $q(x,y)$, which can be approximated by a sample average over the MC data. Therefore, the only challenge lies in the estimation of the function $w(y,x,\theta)$. Recall that $w(y,x,\theta)$ is defined as the ratio of two conditional densities $\frac{p(y|x,\theta)}{q(y|x)}$. Chan and Nachman (2023) proposed to learn $w(y,x,\theta)$ by factorizing it into the product of two density ratios
@@
Let $X_{i}\sim p_{X},\theta_{i}\sim p_{\theta},Y_{i}\sim p(\cdot|X_{i},\theta_{i})$, and $X_{i}^{\prime}\sim q_{X},\theta_{i}^{\prime}\sim q_{\theta},Y_{i}^{\prime}\sim q(\cdot|X_{i})$. Let $f_{1}:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow[0,1]$ be the Bayes optimal classifier distinguishing dataset $\mathcal{D}_{1}=\{X_{i},Y_{i},\theta_{i}\}$ from $\mathcal{D}_{2}=\{X^{\prime}_{i},Y^{\prime}_{i},\theta_{i}^{\prime}\}$. Let $f_{2}:\mathcal{X}\times\Theta\rightarrow[0,1]$ be the Bayes optimal classifier distinguishing dataset $\tilde{\mathcal{D}}_{2}=\{X^{\prime}_{i},\theta^{\prime}_{i}\}$ from $\tilde{\mathcal{D}}_{1}=\{X_{i},\theta_{i}\}$. Then the outputs of the classifiers satisfy:
@@
The procedure outlined in Proposition 3 was also used in Chan and Nachman (2023), although a formal proof was not provided. In this setting, the auxiliary variable $\theta_{i}^{\prime}$ does not influence the synthetic data $X_{i}^{\prime},Y_{i}^{\prime}$; rather, it functions solely as a supporting variable for classifier training. While the distributions of $\theta_{i}$ and $\theta_{i}^{\prime}$ may differ, as allowed by the proposition, in practice, it might be preferable to set $p_{\theta}=q_{\theta}$ to avoid potential numerical instability. A convenient choice is to use a uniform distribution over the parameter space. Similarly, although the proposition specifies $X_{i}\sim p_{X},X_{i}^{\prime}\sim q_{X}$, there is no restriction against using the same distribution for both. Therefore, a practical procedure can be summarized as follows:
@@
    Sample $X_{i}^{\prime},X_{i}$ from the particle-level MC distribution $q_{X}$.
@@ 4.2 Neural network architecture and training
The neural network classifier for estimating density ratios during POF iteration is implemented in TensorFlow and Keras (Abadi et al., 2016; Chollet, 2017). The network contains three hidden layers with 50 nodes per layer and employs the ReLU activation function. The output layer consists of a single node with a sigmoid activation function. Training is performed with the Adam optimizer (Kingma and Ba, 2017) with learning rate $\eta=0.001$ using a weighted binary cross-entropy loss. The model is trained for up to 20 epochs with a batch size of 10,000, and early stopping with a patience of 3 epochs is applied, that is, training stops if the validation loss does not improve for 3 consecutive epochs. None of the hyperparameters were optimized.
Figure 4: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 23 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
The neural network classifiers for estimating the $w$ function ($w(y,x,\theta)=p(y|x,\theta)/q(y|x)$) are implemented in PyTorch (Paszke et al., 2019). [^2] The reason we used two different frameworks is that the original implementation of OmniFold was in TensorFlow/Keras, while the codebase for learning $w$ function from Chan and Nachman (2023) was in PyTorch. There is no technical reason preventing from using a single framework. The same architecture is employed for both classifiers in Proposition 3. The network contains three hidden layers with 50 nodes per layer and employs the ReLU activation function. Batch normalization is applied after each hidden layer, and a dropout layer with a rate of 0.1 is added after the second hidden layer. The output layer consists of a single node with a sigmoid activation function. Training uses the Adam optimizer with learning rate $\eta=0.001$ and the weighted binary cross-entropy loss. This classifier is trained for up to 1000 epochs with a batch size of 10,000, and early stopping with patience 10 is used. In addition, we train an ensemble of 10 networks with bootstrap resampling to reduce the variance of the estimated w function. The range of nuisance parameter $\theta$ used in training is set to be $[0.5,2.0]$.
@@ 4.3 Results
Figure 7: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 56 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
Moreover, Figure 4 shows the evolution of the nuisance parameter $\theta$ and the goodness-of-fit statistic for the step-1 classifier, defined in Eq. (25). The top plot shows that the nuisance parameter converges to the true value within a few iterations, regardless of the initial value. The bottom plot shows the goodness-of-fit statistic converging to 1 for all initial values, indicating that the reweighted distribution $\tilde{q}(y)$ is close to the target distribution $p(y)$. As discussed in Section 3.1, the results shown in Figures 23 correspond to the solution with the highest goodness-of-fit statistic, although in this case study all solutions give equally good fits after a few iterations.
@@ 5.1 Dataset
To demonstrate this setup in practice, we use high-fidelity simulations from the CMS Experiment at the Large Hadron Collider. In particular, the data were generated with Pythia 6.426 (Sjöstrand et al., 2006) using the Z2 tune (Chatrchyan and others, 2011) and interfaced with a Geant4-based (Agostinelli and others, 2003; Allison and others, 2006, 2016) detailed detector simulation of the CMS experiment (Chatrchyan and others, 2008). This dataset comes from the CMS Open Data Portal (CMS collaboration, 2016; CMS Collaboration, 2016a, b) and is processed into an MIT Open Data format (Komiske et al., 2019a, b, c, 2020). We use the events as both “simulation” and “data” in order to have a known target for testing.
The datasets from this collection are sorted by the parton-level hard-scattering scale $\hat{p}_{T}$ from Pythia, which is in general different from the jet-level transverse momentum $p_{T}$ we are interested in studying. For simplicity, we consider one slice of the collection with 600 GeV $\hat{p}_{T}<$ 800 GeV. This slice corresponds to sufficiently high momentum jets that effects from the triggering system are not relevant. Particles (at truth level) or particle flow candidates (at reconstructed/detector level) are used as inputs to jet clustering, implemented using FastJet 3.2.1 (Cacciari et al., 2012; Cacciari and Salam, 2006) and the anti-$k_{t}$ algorithm (Cacciari et al., 2008) with radius parameter $R=0.5$. The total number of events is 43,892. We consider the two highest $p_{T}$ jets at both particle (truth) and detector (reconstructed) levels, denoted as $p_{T,1}^{\text{truth}},p_{T,2}^{\text{truth}},p_{T,1}^{\text{reco}},p_{T,2}^{\text{reco}}$. Then our observables are
@@ 5.3 Results
Moreover, Figure 10 shows that the nuisance parameter does not necessarily converge to a good estimate. In particular, when initialized at $\theta^{(0)}=1.0$ or $1.1$, the estimated value stabilizes around $\hat{\theta}\approx 1.35$. In these cases, the corresponding goodness-of-fit statistic also fails to approach 1, indicating that the reweighted distribution $\tilde{q}(y)$ does not adequately match the observed distribution $p(y)$. This behavior suggests that the algorithm may converge to a local, rather than global, maximum of the likelihood if the starting value $\theta^{(0)}$ is far from the optimal value. The results reported in Figures 89 are obtained by selecting, among all runs, the solution achieving the highest goodness-of-fit statistic, thereby avoiding this issue. Additional experiments with different jet energy resolution are provided in the supplementary material. These experiments yield consistent results: POF reliably recovers the true particle-level distribution, whereas OF fails in the presence of a misspecified nuisance parameter. However, the choice of the initial nuisance parameter remains important for POF, as starting too far from the true value can lead to suboptimal estimates. For both OF and POF, the reweighted detector-level distributions may also fail to perfectly match the experimental data, particularly when the discrepancy between the true $\theta$ and the nominal $\theta$ used in simulation is large. These findings highlight the importance of employing multiple initializations and selecting the best solution based on the goodness-of-fit statistic.
Figure 10: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 89 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@ 6 Discussion
Another important direction for future work is uncertainty quantification of the unfolding results. The current POF algorithm does not provide uncertainty estimates for either the unfolded distribution or the nuisance parameters. Even in the original OmniFold algorithm, it is unclear how to propagate the uncertainty in the classifier-based density ratio estimates to the final unfolded result. One possible (computationally expensive) solution is to use bootstrap resampling (Canelli et al., 2025) to estimate the uncertainty of the unfolded distribution and nuisance parameters. However, rigorous analysis of uncertainty quantification still remains an open problem in this setting.
@@ A.1 Gaussian Data
    Figure 13: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 1112 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 16: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 1415 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 19: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 1718 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 22: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 2021 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@ A.2 CMS Open Data
    Figure 25: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 2324 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 28: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 2627 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 31: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 2930 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.
@@
    Figure 34: Evolution of the nuisance parameter and the step-1 classifier’s goodness-of-fit statistic for the POF algorithm under different initializations $\theta^{(0)}$. The results shown in Figures 3233 correspond to the solution yielding the highest goodness-of-fit statistic. Top: Updated estimates $\hat{\theta}$ across iterations for different initializations $\theta^{(0)}$. Bottom: Goodness-of-fit statistic of the step-1 classifier at each iteration.